# Coupling of phase transition, anharmonicity, and thermal transport in CaSnF6

### 31 Mar 2026 v2

@@ Abstract
Understanding the coupling between structural phase transitions and thermal transport is essential for designing functional materials with tunable properties. Here, we investigate this interplay in CaSnF6 by combining first-principles calculations with a machine-learned neuroevolution potential that enables large-scale molecular dynamics simulations across a wide temperature range. The simulations accurately capture the first-order structural phase transition and associated lattice dynamics. We show that the negative thermal expansion originates from low-energy rigid unit modes involving cooperative rotations of corner-sharing [CaF6]4- octahedra, which induce bond-angle bending and volume contraction. At the same time, strong anharmonicity, dominated by four-phonon scattering, plays a central role in suppressing lattice thermal conductivity ($\kappa_{L}$). Crucially, non-equilibrium simulations reveal a pronounced non-monotonic anomaly in $\kappa_{L}$ near the phase transition, deviating from the conventional $\sim 1/T^{\alpha}$ behavior and providing direct transport evidence of lattice reconstruction. These results establish a unified mechanism linking lattice geometry, anharmonic vibrational dynamics, and thermal transport, and highlight the potential of machine-learned potentials for bridging atomic-scale phase transitions with macroscopic transport properties.
@@ III Results and Discussions
The negative thermal expansion in CaSnF6 is intimately linked to its structural framework and vibrational dynamics. Similar to canonical NTE materials such as ScF3, ReO3, and ZrW2O8dove2016negative , the structure consists of corner-sharing [CaF6]4- octahedra forming a three-dimensional flexible network. This topology supports low-energy rigid unit modes (RUMs), involving cooperative rotations of nearly rigid octahedra. To elucidate the microscopic origin of NTE, Fig. 5(b) shows the phonon dispersion projected with mode-resolved Grüneisen parameters. The three lowest-frequency optical branches ($o_{1}$, $o_{2}$, $o_{3}$) exhibit the largest (negative) Grüneisen parameters, indicating their dominant contribution to NTE. The corresponding vibrational eigenmodes at the $\Gamma$ point are shown in Figs. 5(c-e), where the three modes are nearly degenerate. These modes primarily involve rotational motions of [CaF6]4- octahedra coupled with transverse vibrations of F atoms. Upon heating, the transverse motion of F atoms and the cooperative rotations of [CaF6]4- octahedra (RUMs) induce bending of the Ca-F-Sn bond angle, thereby shortening the Ca-Sn distance while largely preserving the Sn-F bond length, leading to a net volume contraction. This effect is directly illustrated in Fig. 5(a), where the Ca-F-Sn bond angle decreases monotonically with temperature, with a reduction of approximately $1.20^{\circ}$ per 100 K. These modes possess negative Grüneisen parameters and are thermally activated, giving rise to the observed macroscopic negative thermal expansion.
@@ IV Conclusion
In summary, we establish a comprehensive picture of the structural phase transition in CaSnF6 and its strong coupling to thermal transport by integrating first-principles calculations with a high-fidelity NEP. The developed NEP accurately reproduces the underlying energy landscape and enables large-scale molecular dynamics simulations that capture the first-order phase transition, bridging the gap between atomistic accuracy and mesoscopic dynamics. From a microscopic perspective, the NTE is shown to originate from low-energy rigid unit modes associated with cooperative rotations of corner-sharing [CaF6]4- octahedra, which induce bond-angle bending and volume contraction. Concurrently, strong anharmonicity, manifested in enhanced 4ph scattering, plays a dominant role in suppressing $\kappa_{L}$. Crucially, HNEMD simulations reveal a pronounced non-monotonic anomaly in $\kappa_{L}$ in the vicinity of the phase transition, providing direct transport-level evidence of lattice reconstruction. This anomalous behavior reflects the competition between phonon hardening, scattering enhancement, and rapid reorganization of vibrational states across the phase boundary. Overall, this work uncovers a unified mechanism linking lattice geometry, anharmonic lattice dynamics, and thermal transport in CaSnF6. The combined framework of machine-learned potentials and multiscale transport simulations provides a robust and general strategy for investigating phase-transition-driven functional responses in complex materials.

### 14 Apr 2026 v3

@@ Abstract
Understanding the coupling between structural phase transitions and thermal transport is essential for designing functional materials with tunable properties. Here, we investigate this interplay in CaSnF6 by combining first-principles calculations with a machine-learned neuroevolution potential that enables large-scale molecular dynamics simulations across a wide temperature range. The simulations accurately capture the first-order structural phase transition and associated lattice dynamics. We show that the negative thermal expansion originates from low-energy rigid unit modes involving cooperative rotations of corner-sharing [CaF6]4- and [SnF6]2- octahedras, which induce bond-angle bending and volume contraction. At the same time, strong anharmonicity, dominated by four-phonon scattering, plays a central role in suppressing lattice thermal conductivity ($\kappa_{L}$). Crucially, non-equilibrium simulations reveal a pronounced non-monotonic anomaly in $\kappa_{L}$ near the phase transition, deviating from the conventional $\sim 1/T^{\alpha}$ behavior and providing direct transport evidence of lattice reconstruction. These results establish a unified mechanism linking lattice geometry, anharmonic vibrational dynamics, and thermal transport, and highlight the potential of machine-learned potentials for bridging atomic-scale phase transitions with macroscopic transport properties.
@@ III Results and Discussions
The negative thermal expansion in CaSnF6 is intimately linked to its structural framework and vibrational dynamics. Similar to canonical NTE materials such as ScF3, ReO3, and ZrW2O8dove2016negative , the structure consists of corner-sharing [CaF6]4- octahedra forming a three-dimensional flexible network. This topology supports low-energy rigid unit modes (RUMs), involving cooperative rotations of nearly rigid octahedra. To elucidate the microscopic origin of NTE, Fig. 5(b) shows the phonon dispersion projected with mode-resolved Grüneisen parameters. The three lowest-frequency optical branches ($o_{1}$, $o_{2}$, $o_{3}$) exhibit the largest (negative) Grüneisen parameters, indicating their dominant contribution to NTE. The corresponding vibrational eigenmodes at the $\Gamma$ point are shown in Figs. 5(c-e), where the three modes are nearly degenerate. These modes primarily involve rotational motions of [CaF6]4- octahedra coupled with transverse vibrations of F atoms. Rotation of [SnF6]2- was also observed at other high-symmetry points. Upon heating, the transverse motion of F atoms and the cooperative rotations of [CaF6]4- and [SnF6]2- octahedra (RUMs) induce bending of the Ca-F-Sn bond angle, thereby shortening the Ca-Sn distance while largely preserving the Sn-F bond length, leading to a net volume contraction. This effect is directly illustrated in Fig. 5(a), where the Ca-F-Sn bond angle decreases monotonically with temperature, with a reduction of approximately $1.20^{\circ}$ per 100 K. These modes possess negative Grüneisen parameters and are thermally activated, giving rise to the observed macroscopic negative thermal expansion.
@@ IV Conclusion
In summary, we establish a comprehensive picture of the structural phase transition in CaSnF6 and its strong coupling to thermal transport by integrating first-principles calculations with a high-fidelity NEP. The developed NEP accurately reproduces the underlying energy landscape and enables large-scale molecular dynamics simulations that capture the first-order phase transition, bridging the gap between atomistic accuracy and mesoscopic dynamics. From a microscopic perspective, the NTE is shown to originate from low-energy rigid unit modes associated with cooperative rotations of corner-sharing [CaF6]4- and [SnF6]2- octahedra, which induce bond-angle bending and volume contraction. Concurrently, strong anharmonicity, manifested in enhanced 4ph scattering, plays a dominant role in suppressing $\kappa_{L}$. Crucially, HNEMD simulations reveal a pronounced non-monotonic anomaly in $\kappa_{L}$ in the vicinity of the phase transition, providing direct transport-level evidence of lattice reconstruction. This anomalous behavior reflects the competition between phonon hardening, scattering enhancement, and rapid reorganization of vibrational states across the phase boundary. Overall, this work uncovers a unified mechanism linking lattice geometry, anharmonic lattice dynamics, and thermal transport in CaSnF6. The combined framework of machine-learned potentials and multiscale transport simulations provides a robust and general strategy for investigating phase-transition-driven functional responses in complex materials.